suppose the demand curve P=75-4*Q.

What price elasticity of demand when the price is $25?
Answer #1

Answer

converting the demand to normal form

P=75-4Q

4Q=75-P

Q=18.75-0.25P

P=25

Q=18.75-0.25*18=18.75-6.25=12.5

dQ/dP=-0.25 .......... by the first differentiation of the demand curve

E=elastcity =(dQ/dP)* (P/Q)

=-0.25*(25/12.5)

=-0.5

the elastcity is -0.5